ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

20. Special Activities of the ICM 1998

1991 MS Classification: Siebenmann, Laurent, Université Paris-Sud, France A collection of mathematical software on CD-ROM This poster will introduce non-commercial mathematical software appearing on a new CD-ROM for mathematicians – which I have been editing with K. Burdzy and R. S. Palais; see

  http://www.math.washington.edu/~burdzy/MathCD.html,
  http://rsp.math.brandeis.edu/MathCD.html,
  http://topo.math.u-psud.fr/~lcs/MathCD.html.

Individual mathematicians have contributed software representing in all a huge number of years of programming blended with research. Various stategies for assuring the portability and the longevity of academic software will be mentioned, along with questions of archiving, reviewing, distribution and copyright. We can hopefully demonstrate some software with strong visual content using your laptop computer. One is a highly portable setup for viewing 3-dimensional polyhedra stereographically; I am preparing a demo with R. Kenyon.